OpenStudy (anonymous):

In angle ABC, AB=2 and AC=11. Find angle C to the nearest degree.

OpenStudy (anonymous):



OpenStudy (anonymous):

sin(x) =2/11 x=sin^-1(2/11) =10.48?

OpenStudy (anonymous):

use tangent